# Unsourced Random Access: A Comprehensive Survey

### v2

@@ II-B2 Grant-Free Random Access
Grant-free RA is another RA candidate to enable communication of a large number of users where the users transmit their data by employing the shared system resources without any prior grant from the BS, which reduces the latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks, which means that there is user identity during the data transmission phase, and the receiver can utilize this information for activity detection or user separation [36]. This is called sourced (grant-free) random access, which one of solutions based on it, grant-free NOMA, is the prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that, the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection, channel estimation, and data recovery, and synchronization with minimal overhead [37].
Among the two main categories of NOMA, the power-domain NOMA is hard to extend to the grant-free scenario as the successful recovery of the transmitted packets at the receiver heavily depends on the power difference of the users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery [38]. For instance, a grant-free version of SCMA is proposed in [39]. Furthermore, the inherent sparsity of the user activity in grant-free MA is utilized in [40, 41, 42, 43] through compressive sensing algorithms for multiuser detection.
@@ III URA: A Recent Paradigm
To address this problem, Polyanskiy proposed the URA paradigm in [11] where the data transmission and user identification problems are decoupled via the utilization of a common codebook by all users. So, the receiver is not interested in user identities during the data transmission, with the decoding being done only up to a permutation of the transmitted messages.
@@ III-A General System Model
The URA system model in a synchronous setup as follows. Considering a URA system with an $M_{r}$-antenna receiver and single-antenna users, in which $K_{a}$ out of $K_{T}$ users are active at a given time frame of length $n$, the received signal, $\mathbf{Y}\in\mathbb{C}^{M_{r}\times n}$, in the absence of synchronization errors can be written in the following general form [13]
@@
with the following common models in the URA literature
$\begin{cases}M_{r}=1,\mathbf{h}_{i}=1,&\text{GMAC}\\ M_{r}=1,\mathbf{h}_{i}\sim\mathcal{CN}(0,1),&\text{single-antenna fading channel}\\ M_{r}>1,\mathbf{h}_{i}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{r}}),&\text{MIMO fading channel}\\ \end{cases}$ (2)
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the length-$n$ transmitted signal of the $i$-th user corresponding to its message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of length $B$, and ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ is the circularly-symmetric additive white Gaussian noise (AWGN) whose elements are drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$, with $\sigma_{z}^{2}$ being the power of noise. In the literature, it is assumed that each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine a list of transmitted messages, $\mathbf{u}_{i}^{\prime}s,\ \forall i=1,2,...,K_{a}$, by processing the received signal $\mathbf{Y}$.
Figure 3: An illustration of a system where a large number of potential devices (users) try to communicate with an access point (AP), with only a small subset of them being active at each time frame.
In the URA literature, the common setup involves users having messages of fixed length $B$ bits from a set of the possible messages $[M]$ with cardinality $M$, and the objective is to minimize the required energy-per-bit for a target per-user probability of error (PUPE), as the performance metric. The energy-per-bit can be written as
$\dfrac{E_{b}}{N_{0}}=\dfrac{nP}{c\sigma_{z}^{2}B},$ (3)
where $P$ denotes the average power of each user per channel use, i.e., $\norm{\mathbf{x}_{i}}^{2}\leq nP$, and we have $c=2$ for real-valued and $c=1$ for complex channels. Also, PUPE is defined in terms of the probability of false-alarm, $p_{fa}$, and the probability of misdetection, $p_{md}$, as
@@ III-B Challenges
In the following sections, we review the existing URA solutions in the literature. However, in order to have a better understanding of the URA setup and the reasoning behind the provided solutions in the literature, we first need to know the main challenges faced in a URA setup, which will be hinted at in the rest of this section.
Interference: Interference stands out as the primary challenge in MA channels, especially random access due to less coordination. It has an even more detrimental effect on the URA setup since there is no coordination between the receiver and the users. Failure to incorporate strategies for reducing or canceling it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we analyze the treating interference as noise (TIN) scheme, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. The signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy for a GMAC scenario can be written as
@@
[…]
@@
This equation illustrates that when the number of users surpasses a specific threshold (more specifically, when it is much greater than the inverse of transmit signal-to-noise ratio (SNR), i.e., $K\gg\frac{\sigma^{2}_{z}}{P}$, which is not a very large value in many envisioned URA applications), the SINR becomes inversely proportional with $K_{a}$. So, boosting the transmit power does not result in an enhanced SINR or, equivalently, the system performance. Consequently, in systems such as URA, the relatively large number of active users significantly rise in $K_{a}$ (equivalent to an increase in interference) adversely impacts the system performance.
There are two main techniques in the URA literature to alleviate the effects of interference, with the most common one being SIC. SIC is a useful technique for interference mitigation in wireless multiuser scenarios, which removes the effects of successfully detected messages from the received signal to decrease the interference before another round of decoding. Especially in URA, SIC is an essential step in achieving low error rates due to heavy multiuser interference. The other technique for interference management involves randomly separating users into different groups, e.g., separation in time by slotting the time frame, in the power domain by employing different transmit power levels, or in the code domain by multiplying or concatenating a known sequence to the data segment.
@@
    Time-domain separation: Dividing the time frame into multiple slots and letting each user randomly select a slot to transmit its data can lead to enhanced system performance. In a slotted structure, the trade-off is between the effects of decreased interference due to the reduced number of interfering users in each slot and decreased channel coding power due to the fewer channel uses available to each user. Therefore, optimizing the number of slots as a function of system parameters such as active user load plays an important role in benefiting from the slotting.
@@
    Code-domain separation: In some URA schemes, each user randomly selects a preamble/pilot/pattern and uses it to prepare its data for transmission. One way is to transmit a concatenation of a preamble or pilot and the main signal. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. Then, minimum mean square error (LMMSE) or maximum ratio combining (MRC) solutions can be employed to separate the user packets, before decoding them using single-user channel decoders. Another way is to multiply a randomly selected pattern with the main signal, which is used in random-spreading-based solutions. Due to limited available power and time resources, employing these techniques usually decreases the amount of power and time resources dedicated to the data (main signal) part. So, in order for them to be effective in interference management, the choices of type, length, and power of the pilots and pattern sequences are of great importance.
@@
    Power-domain separation: An alternative approach for user separation, especially effective in GMAC scenarios, is the so-called power diversity technique [44], where each user randomly picks a transmit power level from a limited set with a certain probability distribution, satisfying the overall power constraint. Since the users are now separated by different power levels, the receiver, with a high probability, decodes the users with the highest power level first, and after each SIC, it moves to decode the lower power levels. The appropriate design of the power diversity parameters, namely, the number of power levels, their values, and their probability distribution, is very important in achieving the power diversity gain.
Collisions: In many classic random access setups, such as ALOHA, collision is used interchangeably with interference, i.e., when user packs are fully or partially received at the same time at the receiver. However, in the URA literature, collision by default means the collision of either user messages or user pilots, which are when two or more users have exactly the same messages or pilot sequences, respectively.
@@
    Pilot collision: Many existing fading URA solutions employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users independently pick a preamble/pilot from a common codebook, either randomly or based on a part of their message bits. Hence, if the pilot bits of multiple users are the same (called pilot contamination in massive MIMO literature [45]), a collision occurs, and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the pilot collision probability lower than a desired threshold or develop techniques for its mitigation.
Computational Complexity: In order to increase the number of active users supported by a URA system, larger values of parameters such as time frame length and the number of antenna elements (in MIMO systems) are necessary. However, the computational complexity of URA solutions escalates superlinearly with these parameters as well as the number of active users Consequently, solutions for a URA system with a high dimension, temporal (time frame length), spatial (number of receiver antennas), and/or the number of active users, are anticipated to exhibit a high computational complexity. To mitigate this complexity, various strategies are employed in the literature, such as slotted transmissions and coded compressed sensing (CCS). The idea of CCS is to divide the bit sequence into smaller segments, map each segment to a column of a common codebook, and stitch them using a tree code instead of mapping the entire bit sequence to the columns of a $2^{B}\times n$ codebook, which would lead to a prohibitive receiver complexity. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In this section, we summarized the main challenges in a URA setup along with a high-level discussion on the proposed solutions in the literature. In the two subsequent sections, we take a more detailed look at these solutions, with Section IV dedicated to GMAC and Section V to single-antenna (V.A) and MIMO (V.B) fading setups. We take a systematic approach in presenting the URA solutions in these sections: different contributions are grouped based on their solution approach. In each group, the underlying approach is explained with an emphasis on the main work accompanied by a block diagram of its encoder and decoder. The studies are presented in chronological order. To provide an overall view, at the end of each main subsection, a summary of the primary contributions is presented, and a comparative discussion of their performance is provided, accompanied by some numerical results.
@@ IV URA over GMAC
In this section, we present an overview of GMAC URA. To this end, we first present an achievability bound on the performance of the GMAC URA, as presented in Polyanskiy’s original paper [11], and then, review the existing GMAC URA solutions to achieve it.
@@ IV-A Performance Bounds
In [11], Polyanskiy introduces the URA paradigm and obtains an achievability bound on its performance on the following GMAC model with single-antenna users and receiver:
$\mathbf{y}=\mathbf{x}_{1}+\dots+\mathbf{x}_{K_{a}}+\mathbf{z},$ (7)
where $K_{a}$ is the number of active users (i.e., the active user load), $\mathbf{x}_{i}$ is the transmitted signal of the $i$-th user, $\mathbf{y}$ is the received signal and $\mathbf{z}$ is the circularly-symmetric AWGN noise vector with independent and identically distributed (i.i.d.) elements, i.e., $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$.
In [11], an $(M,n,\epsilon)$ RA code, where $\epsilon$ is the target PUPE, is defined as follows
Definition 1.
An $(M,n,\epsilon)$ RA code for the $K_{a}$-user channel $P_{Y\mid X^{K_{a}}}$ is a pair of (possibly randomized) maps, the encoder $f:[M]\rightarrow\mathcal{X}^{n}$ and the decoder $g:\mathcal{Y}^{n}\rightarrow\binom{[M]}{K_{a}}$, satisfying
$\frac{1}{K_{a}}\sum\limits_{j=1}^{K_{a}}\mathbb{P}[E_{j}]\leq\epsilon,$ (8)
where $E_{j}\triangleq\{W_{j}\notin\mathcal{L}\}\cup\{W_{j}=W_{i}\hskip 5.69054pt\text{for some }i\neq j\}$ is the error event for the $j$-th user, $W_{1},\dots,W_{K_{a}}$ are independent and uniform on $[M]$ messages, and $\mathcal{L}$ is the list of the estimated user messages that is produced by the decoder. Note that this error definition consists of the misdetection error and the error resulting from users having the same message sequence. Firstly, it does not take into account the false alarm probability. The reason is that since the number of active users is assumed to be known and a fixed-size decoding list of length $K_{a}$ is used at the receiver in [11], the false alarm probability becomes the same as the misdetection error, so there is no need for its inclusion in the error definition. However, the number of active users is not known by the receiver in practice, and needs to be estimated as is done in other URA solutions, either explicitly or implicitly. Therefore, for a more complete analysis of URA systems, both misdetection and false alarm errors need to be taken into account in the error definition. Another point is that it is assumed in this error definition that any instance of two users having the same message leads to an error, which is not valid. However, this assumption does not violate the achievability of the proposed bound as it assumes a higher error.
The following achievability bound has been used as a benchmark for the energy efficiency of URA over a real-valued AWGN channel.
Theorem 1 (Random coding achievability bound [11]).
Fix $P^{\prime}<P$. There exists an $(M,n,\epsilon)$ random-access code for $K_{a}$-user GMAC satisfying power-constraint $P$ and
$\epsilon\leq\sum\limits_{i=1}^{K_{a}}\frac{t}{K_{a}}min(p_{t},q_{t})+p_{0},$ (9)
where $p_{0}=\frac{\binom{Ka}{2}}{M}+K_{a}\mathbb{P}\Bigg{[}\frac{1}{n}\sum\limits_{j=1}^{n}Z_{j}^{2}>\frac{P}{P^{\prime}}\Bigg{]}$, $p_{t}=e^{-nE(t)}$, $E(t)=\max\limits_{0\leq\rho,\rho_{1}\leq 1}-\rho\rho_{1}tR_{1}-\rho_{1}R_{2}+E_{0}(\rho,\rho_{1})$, $E_{0}=\rho_{1}a+\frac{1}{2}\log(1-2b\rho_{1})$, $a=\frac{\rho}{2}\log(1+2P^{\prime}t\lambda)+\frac{1}{2}\log(1+2P^{\prime}t\mu)$, $b=\rho\lambda-\frac{\mu}{1+2P^{\prime}t\mu}$, $\mu=\frac{\rho\lambda}{1+2P^{\prime}t\lambda}$, $\lambda=\frac{P^{\prime}t-1+\sqrt{D}}{4(1+\rho_{1}\rho)P^{\prime}t}$, $D=(P^{\prime}t-1)^{2}+4P^{\prime}t\frac{1+\rho\rho_{1}}{1+\rho}$, $R_{1}=\frac{1}{n}\log M-\frac{1}{nt}\log(t!)$, $R_{2}=\frac{1}{n}\log\binom{K_{a}}{t}$, and $q_{t}$ is defined as
$q_{t}=\inf_{\gamma}\mathbb{P}[I_{t}\leq\gamma]+exp(n(tR_{1}+R_{2})-\gamma),$ (10)
where the random variable $I_{t}$ is defined as
$I_{t}=\min_{S_{0}}i_{t}(c(S_{0});Y|c(S_{0}^{c}),$ (11)
where $c(S_{0})\triangleq\sum\limits_{j\in S_{0}}c_{j}$ for $c_{1},\dots,c_{M}\sim\mathcal{N}(0,P^{\prime})$ and the minimum is over all $t$- subsets of $[K_{a}]$. The information density $i_{t}$ is defined as
$i_{t}(a;y|b)=nC_{t}+\frac{\log e}{2}\Big{(}\frac{\norm{y-b}_{2}^{2}}{1+P^{\prime}t}-\norm{y-a-b}_{2}^{2}\Big{)},$ (12)
where $C_{t}=\frac{1}{2}\log(1+P^{\prime}t)$.
The idea behind the proof of this theorem is as follows: Let $S=\{W_{1},\dots,W_{K_{a}}\}$ be a random $K_{a}$-subset of $[M]$. We can bound the error event $F_{t}\triangleq\{\lvert S\setminus\hat{S}\rvert=t\}$ with two different methods. The first one is a Gallager-type bound, defining the error events using the generic subsets of size $t$ in $[K_{a}]$ and $[M]\backslash[K_{a}]$, respectively, using the Chernoff bound and applying Gallager’s $\rho$-trick twice. The second one utilizes the information density in (12), considering the event $\tilde{F}=\{I_{t}<\gamma\}$ with the fixed $\gamma$ and $I_{t}$ is given in (11), and taking union over $S_{0}$.
@@
The random coding bound is an information-theoretic performance limit, and it is derived without any complexity constraints as it assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound. In order to illustrate the ineffectiveness of common RA solutions in a URA setup, we first compare their performance with Polyanskiy’s URA achievability result.
In slotted ALOHA, a user packet is successfully recovered if there is no collision and the single-user decoding is successful, as the collided packets are discarded. Moreover, using the normal approximation in [46], the finite blocklength capacity of TIN can be approximated as
$\log M\approx nC_{TIN}(P)-\sqrt{\frac{nP\log^{2}e}{1+K_{a}P}}Q^{-1}(\epsilon),$ (13)
where $C_{TIN}(P)=\frac{1}{2}\log\left(1+\frac{P}{1+(K_{a}-1)P}\right)$, $Q^{-1}(\cdot)$ denotes the inverse Q-function.
@@
In [11], the performances of these approaches are evaluated assuming that there are $K_{a}$ active users transmitting $B=100$ bits of information through a transmission time frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by subsequent GMAC URA works for performance comparison, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
@@
The required $E_{b}/N_{0}$ versus number of active users for these schemes are plotted in Fig. 4 along with the random coding bound derived in [11] and a converse bound for $K_{a}$-GMAC from [47]. The results in Fig. 4 show that the conventional RA techniques (i.e., slotted ALOHA and TIN) are energy inefficient in a URA setup as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. The performance of the grant-based TDMA scheme is also plotted in the figure (referred to as Orthogonal M.A) as a benchmark, in which the time frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame, with its required $E_{b}/N_{0}$ being calculated using the normal approximation in [46]. However, note that the TDMA scheme is not feasible in a URA setup, mainly due to the lack of coordination. Putting the coordination issue aside, the system works in a regime of $n<K_{a}B$ in some URA applications, which renders orthogonal solutions infeasible.
@@
The random coding bound in [11] is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors of [48] extend this bound to the case that the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [49]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [50], where the alarm message needs to be decoded with a lower error rate than the standard ones. This result is extended to the case of asynchronous GMAC in [51], where the delay in transmission is assumed to be up to a small portion of the time frame. However, the provided achievability bound is a loose bound for the asynchronous case as the authors have employed a fixed time frame decoder in which all the packets that have some overlaps with the end of the time frame are deemed undecodable, leading to huge performance degradation, which is not valid in a fully asynchronous setup, where the time horizon is unbounded. Furthermore, the problem of joint delay and activity detection in the asynchronous setup is investigated in [52] using methods of tropical geometry, which is the study of the geometric properties of polynomials when addition is replaced by minimization and multiplication by ordinary addition. Moreover, the conditions for path delays for zero-error detection are identified.

### v3

@@ II-B2 Grant-Free Random Access
Grant-free RA is another RA candidate that enables communication among a large number of users, where users transmit their data by utilizing shared system resources without requiring a prior grant from the BS, thereby reducing latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks, which means that there is user identity during the data transmission phase, and the receiver can use this information for activity detection or user separation [36]. This is also called sourced random access. Among the grant-free RA solutions, grant-free NOMA is a prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection, channel estimation, data recovery, and synchronization with minimal overhead [37].
Among the two main categories of NOMA, power-domain NOMA is challenging to extend to the grant-free scenario, as the successful recovery of transmitted packets at the receiver heavily depends on the power difference between users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery [38]. For instance, a grant-free version of SCMA is proposed in [39]. Furthermore, the inherent sparsity of the user activity in grant-free MA is utilized in [40, 41, 42, 43] through compressive sensing algorithms for multiuser detection.
@@ III URA: A Recent Paradigm
To address this problem, Polyanskiy proposed the URA paradigm in [11] where the data transmission and user identification problems are decoupled via the utilization of a common codebook by all users. Therefore, the receiver is not interested in user identities during data transmission, with decoding being performed only up to a permutation of the transmitted messages.
@@ III-A General System Model
For brevity, we present the general system model for a URA system with a multiple-antenna receiver and complex channel coefficient, since conversions to real channel coefficients and a single-antenna receiver can easily be obtained as special cases. We also assume a quasi-static channel, where the channel coefficients remain constant over the transmission frame  [^1] The transmission frame is the collection of available resources in different dimensions, such as time and frequency, to be used for transmission. We call it frame for brevity hereafter., as the setup can easily be extended to cases where the channel coefficients vary over the frame by replacing the frame with a coherence time block or a time slot over which the channel remains constant. Note that these assumptions are made for the sake of brevity in presenting the general URA system model; however, we have investigated all existing URA solutions in the literature, regardless of their specific system models. Considering a URA system with an $M_{r}$-antenna receiver and single-antenna users, in which $K_{a}$ out of $K_{T}$ users are active at a given frame of length $n$ [11], the received signal, $\mathbf{Y}\in\mathbb{C}^{M_{r}\times n}$, in the absence of synchronization errors can be written in the following general form [13]
@@
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the $1\times n$ transmitted signal of the $i$-th user corresponding to its message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of length $B$, ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ is the circularly-symmetric additive white Gaussian noise (AWGN) with independent and identically distributed (i.i.d.) elements drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$, with $\sigma_{z}^{2}$ being the noise power, and ${\mathbf{h}}_{i}\in\mathbb{C}^{M_{r}\times 1}$ is the vector of channel coefficients. Note that in the URA literature over GMAC, a single-antenna receiver and unit channel coefficients are assumed. However, the default assumption for fading channels is quasi-static fading; i.e., the channel coefficients remain constant across a frame and change from one frame to the next.
Figure 4: An illustration of a system where a large number of potential users (devices) try to communicate with a receiver (access point), with only a small subset of them being active at each time frame.
In the URA literature, it is assumed that each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine the list of transmitted messages, $\mathbf{u}_{i}$’s, $\forall i=1,2,...,K_{a}$, by processing the received signal $\mathbf{Y}$ so that the required energy-per-bit for a target per-user probability of error (PUPE), as the performance metric, is minimized [57, 12, 102]. The PUPE captures both missed detections and false alarms. Specifically, it is defined as the sum of probabilities that an active user’s message is not correctly included in the decoded list (probability of misdetection, $p_{md}$), and that an invalid message appears in the list (probability of false-alarm, $p_{fa}$); i.e., [14]
@@
The energy per bit can be written as
$\dfrac{E_{b}}{N_{0}}=\dfrac{nP}{c\sigma_{z}^{2}B},$ (3)
where $P$ denotes the average power of each user per channel use, i.e., $\norm{\mathbf{x}_{i}}^{2}\leq nP$, and we have $c=2$ for real-valued and $c=1$ for complex-valued channels.
@@ III-B Challenges
In the following sections, we review existing URA solutions in the literature. However, to gain a better understanding of the URA setup and the reasoning behind the solutions provided in the literature, we first need to identify the main challenges faced in a URA setup, which will be discussed in the remainder of this section.
Interference: Interference stands out as the primary challenge in MA channels, especially in random access setups such as URA, due to the lack of coordination between the receiver and the users. Failure to incorporate strategies to reduce or cancel it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we analyze the treating interference as noise (TIN) scheme, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. The signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy for a GMAC scenario can be written as
@@
where $P$ is the transmit power of each user. Note that the SINR in a fading scenario is approximately the same as in (1) since we can approximate $\frac{1}{M_{r}}\|\mathbf{h}_{i}\|^{2}\approx 1,\ \forall i$, where $\|\cdot\|$ denotes the Euclidean norm. For a sufficiently large active user load $K_{a}$, the SINR can be written as
@@
This equation illustrates that when the number of users exceeds a specific threshold (more specifically, when it is much greater than the inverse of the signal-to-noise ratio (SNR), i.e. $K\gg\frac{\sigma^{2}_{z}}{P}$, which is not a very large value in many envisioned URA applications), the SINR becomes inversely proportional to $K_{a}$. Therefore, boosting the transmit power does not result in an enhanced SINR, or equivalently, improved system performance. Consequently, in systems such as URA, the relatively large number of active users has an adverse impact on system performance.
There are two main techniques in the URA literature to alleviate the effects of interference, with the most common being SIC, which removes the effects of successfully detected messages from the received signal to decrease the interference before another round of decoding. Especially in URA, SIC is a crucial step in achieving low error rates due to heavy multiuser interference. Other techniques for interference management involve randomly separating users into different groups, such as slotting the time frame, employing different transmit power levels, and multiplying or concatenating a known sequence to the data segment.
@@
    Time-domain separation: Dividing the frame into multiple slots and letting each user randomly select a slot to transmit its data can lead to improved system performance. In a slotted structure, the level of interference is reduced due to the lower number of interfering users in each slot. At the same time, the gain of channel coding is also decreased because fewer channel uses are available to each user. Thus, optimizing the number of slots based on system parameters, such as active user load, is crucial for maximizing the advantages of slotting.
@@
    Code-domain separation: In some URA schemes, each user randomly selects a preamble/pilot/pattern and uses it to prepare its data for transmission. One way is to transmit a concatenation of a preamble or pilot and the main signal. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. Then, minimum mean square error (MMSE) or maximum ratio combining (MRC) solutions can be employed to separate the user packets, before decoding them using single-user channel decoders. Another way is to multiply a randomly selected pattern by the main signal, which is used in random-spreading-based solutions. However, employing these techniques usually decreases the amount of power and time resources dedicated to the data (main signal) part. Therefore, for interference management to be effective, the choices of type, length, and power of the pilots and pattern sequences are crucial.
@@
    Power-domain separation: An alternative approach for user separation, especially effective in GMAC scenarios, is the so-called power diversity technique [44], where each user randomly picks a transmit power level from a limited set with a specific probability distribution, satisfying the overall power constraint. Since the users are now separated by different power levels, the receiver, with a high probability, decodes the users with the highest power level first, and after each SIC step, it moves to decode the lower power levels. The appropriate design of the power diversity parameters, namely, the number of power levels, their values, and their probability distribution, is crucial for optimizing the system performance.
Collisions: In many classical random access setups, such as ALOHA, collision is used interchangeably with interference, i.e., when user packets are fully or partially received at the same time at the receiver. However, in the URA literature, collision by default refers to the collision of either user messages or user pilots, which occurs when two or more users have the same messages or pilot sequences, respectively.
@@
    Pilot collision: Many existing works taking into account channel fading employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users independently pick a preamble/pilot from a common codebook, either randomly or based on a part of their message bits. Hence, if the pilot bits of multiple users are the same (called pilot contamination in massive MIMO literature [45]), a collision occurs, and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the pilot collision probability lower than a desired threshold or develop techniques for its mitigation.
Computational Complexity: In order to increase the number of active users supported by a URA system, larger values of parameters such as frame length and the number of antenna elements (in MIMO systems) are necessary. However, the computational complexity of URA solutions typically increases superlinearly with these parameters, as well as the number of active users. As a result, solutions for a URA system with high dimensions, including temporal (frame length), spatial (number of receiver antennas), and/or the number of active users, are expected to experience high computational complexity. To mitigate this complexity, various strategies are employed in the literature, such as slotted transmissions and coded compressed sensing (CCS). The idea of CCS is to divide the bit sequence into smaller segments, map each segment to a column of a common codebook, and stitch them together using a tree code, rather than mapping the entire bit sequence to the columns of a $2^{B}\times n$ codebook, which would lead to prohibitive receiver complexity. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In this section, we summarized the main challenges in a URA setup along with a high-level discussion on the proposed solutions in the literature. In the two subsequent sections, we provide a more detailed examination of these solutions, with Section IV dedicated to GMAC and Section V covering single-antenna (V.A) and MIMO (V.B) fading setups. We take a systematic approach in presenting the URA solutions in these sections: different contributions are grouped based on their solution approach. In each group, the underlying approach is explained with an emphasis on the main structure of its encoding and decoding approaches, and the relevant works are presented in chronological order. To provide an overall view, at the end of each main subsection, a summary of the primary contributions is presented, and a comparative discussion of their performance is given, accompanied by some numerical results.
@@ IV URA over GMAC
In this section, we provide an overview of URA works over GMAC. To this end, we first describe an achievability bound on the performance of the URA over GMAC, as developed in Polyanskiy’s original paper [11], and then review the existing URA over GMAC solutions to achieve it.
@@ IV-A Performance Bounds
Polyanskiy provides a random coding achievability bound on the performance of URA over GMAC in [11], as a combination of a Chernoff-type bound and an information density-based one, which has been used as a benchmark for the energy efficiency in the URA over GMAC literature. The random coding bound is an information-theoretic performance limit, derived without any complexity constraints. It assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound.
@@
In [11], the performances of common RA solutions are evaluated in a URA setup, assuming that there are $K_{a}$ active users transmitting $B=100$ bits of information through a transmission frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by subsequent URA works over GMAC for performance comparison, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
@@
The curves for the required $E_{b}/N_{0}$ versus the number of active users for slotted ALOHA, TIN, and orthogonal MA are plotted in Fig. 5 along with the random coding bound derived in [11] and a converse bound for $K_{a}$-GMAC from [47]. The results in Fig. 5 show that the conventional RA techniques (i.e., slotted ALOHA and TIN) are energy inefficient in a URA setup, as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. For slotted ALOHA, it is assumed that a user packet is successfully recovered if there is no collision and the single-user decoding is successful, while the finite blocklength capacity of TIN is estimated using the normal approximation in [46]. The performance of the grant-based TDMA scheme is also plotted in the figure (referred to as Orthogonal M.A) as a benchmark, in which the frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame, with its required $E_{b}/N_{0}$ being calculated using the normal approximation in [46]. However, note that the TDMA scheme is not feasible in a URA setup, mainly due to the lack of coordination. Leaving the coordination issue aside, the system works in the regime of $n<K_{a}B$ in some URA applications, rendering orthogonal solutions such as TDMA infeasible.
@@
The random coding bound in [11] is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors in [48] extend this bound to the case where the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [49]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [50], where the alarm message needs to be decoded with a lower error rate than the standard ones. An achievability bound for the case of asynchronous GMAC is obtained in [51], where the delay in transmission is assumed to be up to a small portion of the time frame. However, the provided achievability bound is loose for the asynchronous case as the authors have employed a fixed frame decoder in which all the packets that have some overlaps with the end of the frame are deemed undecodable, leading to huge performance degradation, which is not valid in a fully asynchronous setup, where the time horizon is unbounded. The drawbacks in [51] are alleviated in [52], where the authors study a wrap-around version of the model in [51], that is, the last $D_{m}$ symbols are added to the first $D_{m}$ symbols, where $D_{m}$ is the maximum delay. As a result, the receiver observes the superposition of $K_{a}$ cyclically shifted codewords with Gaussian noise, where the noise power in the first $D_{m}$ symbols is doubled. In this way, the energy efficiency of the bound in [51] is significantly improved. The problem of joint delay and activity detection in the asynchronous setup is investigated in [53] using methods of tropical geometry, which is the study of the geometric properties of polynomials when addition is replaced by minimization and multiplication by ordinary addition. Moreover, the conditions for path delays for zero-error detection are identified. Furthermore, the authors in [54] examine a two-user binary adder channel in the asynchronous scenario, and propose an algorithm to avoid the small weight stopping sets in the joint factor graph, which mitigates the error floor due to these sets.